Assume that you have the same volume of aluminum and 80 g of an unknown metal. What is the density of the unknown metal if the Aluminum has a mass of 19 g and a density of 2.7 g/cm3? Can you determine which type of metal it is?
Answer #1

Given Volume of aluminum = volume of unknown metal

Volume of aluminum= mass/density = 19/2.7 cm3

Density of unknown metal = mass/volume = 80/(19/2.7) =11.368 g/cm3

It may be LEAD as its approximate density = 11.34 g/cm3